OpenStudy (anonymous):

Al saves pennies. He agreed to give 5/9 0f his pennies to Bev if she gives 5/9 0f what she got from Al to Carl. In turn Carl will give 5/9 of what he got from Bev to Dani. After that Dani has 875 pennies total. How many pennies does Al start with?

OpenStudy (anonymous):

Al = a. Your starting amount and what you are solving for. Bev (b) receives 5/9a Carl (c) receives 5/9b Dani receives 5/9c Dani = 875 Set up your equation: Carl is \[\frac{ 5 }{ 9 }\left( \frac{ 5 }{ 9 }a \right)\] Dani is 5/9 carl \[\frac{ 5 }{ 9 }\left( \frac{ 5 }{ 9 }\left( \frac{ 5 }{ 9 }a \right) \right)\] Since we know Dani has 875, set that last equation equal to 875 and solve.

OpenStudy (anonymous):

How do u set the last equation to 875?

OpenStudy (anonymous):

\[875 = \frac{ 5 }{ 9 }\left( \frac{ 5 }{ 9 }\left( \frac{ 5 }{ 9 }a \right) \right)\]

OpenStudy (anonymous):

multiply 5/9 * 5/9 * 5/9=875